OpenStudy (anonymous):

\[\tan^{-1} z\] inverses of trig functions

OpenStudy (anonymous):

what u need to know about this function?

OpenStudy (anonymous):

What is your question?

OpenStudy (anonymous):

w(z) = tan−1z show that it is = -i/2*log(i-z)/(i+z)

OpenStudy (anonymous):

@sami-21 @eliassaab

OpenStudy (dumbcow):

http://www.wolframalpha.com/input/?i=arctan%28a%2Bbi%29 looks like it comes from definition of inverse hyperbolic tangent

OpenStudy (anonymous):

ok lets do this we can write it as \[z=\tan(w)\] \[\tan(w)=\frac{\sin(w)}{\cos(w)} \] and \[\sin(w)=\frac{1}{i}(e^{iw}-e^{-iw})\] also \[\cos(w)=(e^{iw}+e^{-iw})\] so it becomes

OpenStudy (anonymous):

@SkykhanFalcon are u fallowing the steps?

OpenStudy (anonymous):

can you solve it for ??

OpenStudy (anonymous):

can u solve this for \[e^{2iw}\] ?

OpenStudy (anonymous):

sorry i wasnt here :/

OpenStudy (anonymous):

ok look at the above can you solve this for \[e^{2iw}\]

OpenStudy (anonymous):

okey thanks

OpenStudy (anonymous):

what did u get after solving?

OpenStudy (anonymous):

just take log of both sides of the above and you will get your result :)

OpenStudy (anonymous):

@sami-21 dude i just regonized your solution is full of wrongs

OpenStudy (anonymous):

mistakes

OpenStudy (anonymous):

@SkykhanFalcon just find one mistake. and let me know

OpenStudy (anonymous):

i was asking arctanz

OpenStudy (anonymous):

the answer \[(-i/2)*\log(i-z)/(i+z)\]

OpenStudy (experimentx):